A bullet of mass 20g is found to pass two points 30m apart in a time interval of 4sec. Calculate kinetic energy of the bullet,if it moves with a constant speed

Grade:12th

2 Answers

Arun
6 years ago
 
velocity of bullet = 30/4 = 15/2
k.e = 20 *15*15 /(2*2*1000) =1.125 j 
where k.e is kinetic energy
j is joules 
 
velocity = distance / time 
Ruchika gupta
4 years ago
V=s/t
V=30/4 = 7.5
Kinetic energy =1/2mv^2
Kinetic energy =1/2×0.02×7.5×7.5
Kinetic energy = 0.5625 joule